a stone is dropped from a cliff onto a flowing river below. if it takes the stone 0.185 minutes to hit the water, how high is the cliff?
1
Expert's answer
2011-02-17T05:53:49-0500
According to the equations of the motion,
h = h0 + v0t - gt2/2
In our case h =0, v0= 0, t = 0.185[minutes] *60 = 11.1 [s], g= 9.8 [m/s2]
thus
h0 = gt2/2 = 9.8* 11.12/2 = 603.729 [m]